OpenStudy (anonymous):

Sabrina is fencing her yard, which is shaped as a rhombus. She used (6x – 7) meters of fencing material for one side and (3x+ 5) meters for the adjacent side. What is the perimeter of the yard?

OpenStudy (lollylau):

Rhombus have four sides of its lengths equal. therefore, 6x-7=3x+5. 9x=12 x=1+1/3 Hope that helped!

OpenStudy (lollylau):

you can calculate the perimeter, can you?